OpenStudy (jhurtado2016):

The area (A) of a circle with a radius of r is given by the fo rmula and its diameter (d) is given by . Arrange the equations in the correct sequence to rewrite the formula for diameter in terms of the area of the circle.

OpenStudy (photon336):

\[A = \pi*r^{2}\] \[Diameter = 2*r \]

OpenStudy (photon336):

@jhurtado2016

OpenStudy (photon336):

the variable in both equations is r \[\sqrt{\frac{ A }{ \pi }} = r \] now \[D = r*2 \] what do you think you can do now?